# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v3

@@ 4 Energy-like Metrics: What Can Go Wrong
One important limitation of these metrics is that they indicate oversmoothing only in the limit of infinitely many layers, when their values converge exactly to zero. Since they measure a form of absolute distance, a small but nonzero value does not provide any meaningful information. On the other hand, convergence to zero corresponds to either perfect alignment with the dominant eigenspace or the collapse of the feature representation matrix to the all-zero matrix. While the former is a symptom of oversmoothing, the latter does not necessarily imply oversmoothing. Moreover, this convergence property requires the weights to be bounded. However, in practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with arbitrarily large (or small) weight magnitudes. This aligns with our intuition and what occurs in the linear case. Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ is dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph. More precisely, the following theorem holds:
@@
[…]
@@
In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely $X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{\top}+R(l)\Big{)}$ for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$. This results in weakly expressive feature representations, independently of the magnitude of the feature weights. This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it, this is e.g. the case when $\lambda_{1}>1$ having $X^{(l)}\rightarrow\infty$.
Another important limitation of these metrics is their dependence on a specific, known dominant eigenspace, which must either be explicitly known or computed in advance. Consequently, their applicability is strongly tied to the specific architecture of the network. In particular, the dominant eigenvector $u$ of $A^{(l)}$ must be known and remain the same for all $l$. This requirement excludes their use in most cases where $A^{(l)}$ varies with $l$, such as when $A^{(l)}$ is the standard weighted adjacency matrix of the graph.
@@ 5 The Rank as a Measure of Oversmoothing
Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing [14]. However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure. A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank [31, 32, 1]. The stable rank is defined as $\text{StableRank}(X)=\|X\|_{*}^{2}/\|X\|_{F}^{2}$, where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm. The numerical rank is given by $\text{NumRank}(X)=\|X\|_{F}^{2}/\|X\|_{2}^{2}$. Finally, given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, the effective rank is defined as
$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$
where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown in Section 6.
In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps addressing the limitations of Dirichlet-like measures. Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one. A small value of the effective rank directly implies that the feature representations are low-rank, suggesting a potentially suboptimal network architecture.
In Figure 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned. This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one. In contrast, continuous rank functions provide a more reliable measure of approximate feature alignment. Later, in Figure 2, we demonstrate that the same phenomenon occurs in GNNs trained on real datasets, where exact feature alignment is rare. In such cases, classical metrics remain roughly constant, whereas the rank decreases, coinciding with a sharp drop in GNN accuracy.
@@
Figure 1: Toy scenarios depicting the behaviour of oversmoothing metrics. Each plot contains 50 nodes, each with two features plotted on the x-y axis. The features are: # 1 of the same value; # 2 perfectly aligned with the same vector; # 3 aligned to the same vector except for one (red) point; # 4 sampled from a uniform distribution. MAD (Sec. 6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 although features are oversmoothing by definition. $E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing. To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$, the first feature was considered in place of $u$ in (1) and (2).
@@ 5.1 Theoretical Analysis of Rank Decay
In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entry-wise nonnegative. While this is a somewhat restrictive setting, our result is the first theoretical proof that oversmoothing occurs independently of the weight (and thus feature) magnitude.
@@
$\text{NumRank}(X)=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$ (4)
The above inequality, together with Theorem 4.1, allows us to establish the convergence of the numerical rank for linear networks.
@@
Consider a linear GCN of the form $X^{(l+1)}=AX^{(l)}W^{(l)}$, where $A$ has a simple dominant eigenvalue $\lambda_{1}$ satisfying $|\lambda_{1}|\geq|\lambda_{2}|$. We have already noted that $\|X\|_{2}\geq\|\mathcal{P}X\|_{2}$, meaning that the numerical rank converges to one if $\|(I-\mathcal{P})X\|_{F}/\|X\|_{2}$ decays to zero. This occurs whenever the features grow faster in the direction of the dominant eigenvector than in any other direction. As established in Theorem 4.1, this is almost surely the case in linear GNNs. As a direct consequence, we obtain the following result:
@@
Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Under the same assumptions as in Theorem 4.1, the following identity holds almost surely:
@@
Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial. In the next subsection, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
5.1.1 The Nonlinear Case
To study the case of networks with nonlinear activations, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to [18, 12] and the reference therein for further details.
Assume all the intermediate features of the network to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$. On $\mathcal{K}$, it is possible to introduce the partial ordering
$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$
where $\bar{\mathcal{K}}$ denotes the closed cone of nonnegative vectors. Given two points $x,y\in\mathcal{K}$ we write $M(x/y)=\max_{i}x_{i}/y_{i}$ and $m(x/y)=\min_{i}x_{i}/y_{i}$. Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$, $d_{H}(x,y)=\log(M(x/y)/m(x/y)).$ Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$. However it is a distance up to scaling; that is, it becomes a concrete distance whenever we restrict ourselves to a slice of the cone. It is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
$d_{H}(Ax,Ay)\leq\beta d_{H}(x,y)\qquad\forall x,y\in\mathcal{K}$ (5)
with $0\leq\beta\leq 1$. The last inequality is a consequence of the fact that $A$ preserves the ordering induced by the cone, i.e. if $x\geq_{\mathcal{K}}y$, then $Ax\geq_{\mathcal{K}}Ay$. In particular, we recall that (5) always hold for $\beta<1$ when $A$ is entry-wise strictly positive, as positive matrices map the whole $\bar{\mathcal{K}}$ in its interior, i.e. $x\geq_{\mathcal{K}}y$ implies $Ax>Ay$. In the next definition, we include all the nonnegative matrices that have the same behaviour as the positive matrices but only on a neighbourhood of a positive vector in $\mathcal{K}$. To this end, we recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
Definition 5.2.
A family of nonnegative irreducible matrices $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$
Note that the above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
Consider now first the case of a linear function that can be represented in the form $F(X)=AXW$, where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively. Then, under mild assumptions, we can prove that the columns of $F(X)$ are closer to the dominant eigenvector $u$ of $A$ as compared to the columns of $X$.
@@ 5.1.1 The Nonlinear Case
Let $F(X)=AXW$ with $A$ nonnegative and irreducible with dominant eigenvector $u\in\mathcal{K}$. Assume also $X$ to be strictly positive and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$. If $A$ is contractive with respect to $u$ and $C=\max_{i}d_{H}(X_{i},u)$ then
$\max_{i}d_{H}(F(X)_{:,i},u)\leq\beta_{C}\max_{i}d_{H}(X_{:,i},u),$
where $Y_{:,i}$ denotes the $i$-th column of $Y$, and $\beta_{C}<1$.
Nonlinear Perron-Frobenius theory extends some of the results about nonnegative matrices to particular classes of nonlinear functions. A function $\sigma\in C(\mathcal{K},\mathcal{K})$ is order preserving if given any $x,y\in\mathcal{K}$ with $x\geq_{\mathcal{K}}y$ then $\sigma(x)\geq_{\mathcal{K}}\sigma(y)$. Moreover $\sigma$ is subhomogenenous if $\sigma(\lambda x)\leq_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and any $\lambda>1$. In particular, it is strictly subhomogeneous if $\sigma(\lambda x)\ll_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda>1$ and homogeneous if $\sigma(\lambda x)=\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda\geq 0$
Subhomogeneity is a useful property and of practical utility. In fact, as discussed in  [38, 27], a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$. For this family of activation functions, Lemma 5.3 combined with arguments from nonlinear Perron–Frobenius theory yield the following main result
@@
Consider a GNN of the form $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(l)}_{:,i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative, for any $l$. Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(l)}$ and that it is also an eigenvector of $\sigma$. Then, if the matrices $\{A^{(l)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}W^{(l)}_{ij}>0$ for any $l$, it holds
@@
We conclude with a formal investigation of the eigenpairs of activation functions that are entry-wise subhomogeneous. Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$. Then one can easily show that $\sigma$ is itself subhomogeneous on +$\mathbb{R}^{N}_{+}$. We have,
Proposition 5.5.
Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R}_{+},\mathbb{R}_{+})$ be order preserving. Then: 1) If $\sigma$ is homogeneous, any positive vector is an eigenvector of $\sigma$. 2) If $\sigma$ is strictly subhomogeneous, the only eigenvector of $\sigma$ in $\mathcal{K}$ is the constant vector.
As a consequence of the above result, we find that the numerical rank of the features collapses to one both for GCNs with LeakyReLU activation function (the setting of [4]) as well as for GATs with any kind of subhomogeneous activation function (the setting of [45]), as the constant vector is always an eigenvector of an entry-wise nonlinear map. For GCNs with any other subhomogeneous activation functions, the convergence of the numerical rank to 1 as stated in Theorem 5.1 may not hold as discussed in Appendix D.
@@
Table 1: Correlation between the classification accuracy and the logarithm of metric values on GNNs with LeakyReLU and of depths from 2 to 24 layers, separately trained on different homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings) and large-scale (OGB-Arxiv) datasets. For Erank and NumRank, we subtract 1 so that both metrics approach zero. The rightmost column reports the ratio of classification accuracy between GNNs with 2 and 24 layers. Some heterophilic datasets may be more resilient to the increasing network depth, in-line with observations from the literature, e.g. [13]. Additional results on other datasets and activation functions are presented in Appendix E, and with additional network components in Appendix F.

### v4

@@ 4 Energy-like Metrics: What Can Go Wrong
One important limitation of these metrics is that they indicate oversmoothing only in the limit of infinitely many layers, when their values converge exactly to zero. Since they measure a form of absolute distance, a small but nonzero value does not provide any meaningful information. On the other hand, convergence to zero corresponds to either perfect alignment with the dominant eigenspace or the collapse of the feature representation matrix to the all-zero matrix. While the former is a symptom of oversmoothing, the latter does not necessarily imply oversmoothing. Moreover, this convergence property requires the weights to be strongly bounded. However, in most practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with weight magnitudes arbitrarily larger than what is prescribed by Cai and Wang (2020); Wu et al. (2023) or Theorem˜3.2. This aligns with our intuition and what occurs in the linear case. Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ is dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph. More precisely, the following theorem holds:
@@
Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Let $\lambda_{1},\lambda_{2}$ be the largest and second-largest eigenvalues (in modulus) of $A$, respectively. Assume the weights $\{W^{(l)}\}_{l=1}^{\infty}$ are randomly sampled from i.i.d. random variables with distribution $\nu$ such that $\int\log^{+}(\|W\|)d\nu+\int\log^{+}(\|W^{-1}\|)d\nu<\infty$, with $\log^{+}(t)=\max\{\log(t),\,0\}$. If $|\lambda_{2}/\lambda_{1}|<1$, then almost surely
@@
In particular, the theorem above implies that $X^{(l)}=\lambda_{1}^{l}\big(uv^{\top}+R(l)\big)$ for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus, when the spectral gap is large $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, even for moderate values of $l$. This results in weakly expressive feature representations, independently of the magnitude of the feature weights. This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it, as it would, for example, be the case when $\lambda_{1}>1$, having $X^{(l)}\rightarrow\infty$.
Another important limitation of Dirichlet-like metrics is their dependence on a specific dominant eigenspace, which must either be explicitly known or computed in advance. Consequently, their applicability is strongly tied to the specific architecture of the network. In particular, the dominant eigenvector $u$ of $A^{(l)}$ must be known and remain the same for all $l$. This requirement excludes their use in cases where $A^{(l)}$ varies with $l$.
@@ 5 The Rank as a Measure of Oversmoothing
Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing Guo et al. (2023b); Roth and Liebig (2023). However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure. A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank royEffectiveRankMeasure2007a; rudelsonSamplingLargeMatrices2006; Arora et al. (2019), defined as follows
$\text{StabRank}(X)=\frac{\|X\|_{*}^{2}}{\|X\|_{F}^{2}},\quad\text{NumRank}(X)=\frac{\|X\|_{F}^{2}}{\|X\|_{2}^{2}},\quad\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right)$
where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm, and given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is defined as the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown in Section 6.
In practice, measuring oversmoothing in terms of a continuous approximation of the rank helps to address the limitations of Dirichlet-like measures. Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one. A small value of the effective rank directly implies that the feature representations are low-rank, suggesting a potentially suboptimal network architecture.
In Figure˜1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned. This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one. In contrast, continuous rank functions provide a more reliable measure of approximate feature alignment. Later, in Figure 2, we demonstrate that the same phenomenon occurs in GNNs trained on real datasets, where exact feature alignment is rare. In such cases, classical metrics remain roughly constant, whereas the rank decreases, coinciding with a sharp drop in GNN accuracy.
@@
Figure 1: Toy scenarios depicting the behaviour of oversmoothing metrics. Each plot contains 50 nodes, each with two features plotted on the x-y axis. The features are: # 1 of the same value; # 2 perfectly aligned with the same vector; # 3 aligned to the same vector except for one (red) point; # 4 sampled from a uniform distribution. MAD (Sec. 6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 although features are oversmoothing by definition. $E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing. $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ where computed using the first feature in place of $u$ in (1) and (2).
@@ 5.1 Theoretical Analysis of Rank Decay
In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption of linear models or nonlinear models with weight matrices that are entry-wise nonnegative. Our results rigorously show that in these settings, oversmoothing can occur independently of the weight (and thus feature) magnitude and shed light on the possible causes of rank decay.
@@
$\text{NumRank}(X)=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}.$ (4)
The above inequality, together with Theorem˜4.1, allows us to establish the convergence of the numerical rank for linear networks.
@@
Consider a linear GCN of the form $X^{(l+1)}=AX^{(l)}W^{(l)}$, where $A$ has a simple dominant eigenvalue $\lambda_{1}$ satisfying $|\lambda_{1}|\geq|\lambda_{2}|$. We have already noted that $\|X\|_{2}\geq\|\mathcal{P}X\|_{2}$, meaning that the numerical rank converges to one if $\|(I-\mathcal{P})X\|_{F}/\|X\|_{2}$ decays to zero. This occurs whenever the features grow faster in the direction of the dominant eigenvector than in any other direction. As established in Theorem˜4.1, this is almost surely the case in linear GNNs. As a direct consequence, we obtain the following result:
@@
Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Under the same assumptions as in Theorem˜4.1, the following identity holds almost surely:
@@
Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial. However, a simplified setting to study is the one where the weights are nonnegative. Indeed, exactly as in the linear case, while generally the rank decreases only on average, considering nonnegative weights yields a monotone decrease.
The Nonnegative Nonlinear Case
To study the case of networks with nonlinear activations, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to lemmensNonlinearPerronFrobeniusTheory2012; gautier2013nonlinear and the reference therein for further details.
We assume all the intermediate features of the network to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$. Here, we consider the partial ordering $x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)$ if and only if $y-x\geq 0\>(y-x>0)$, where the inequalities have to be understood entrywise. Given two points $x,y\in\mathcal{K}$, let
$d_{H}(x,y)=\log\Big(\max_{i}\frac{x_{i}}{y_{i}}\cdot\max_{i}\frac{y_{i}}{x_{i}}\Big)$ (5)
denote the Hilbert distance. Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$. However, it is a distance up to scaling; that is, it becomes a distance whenever we restrict ourselves to a slice of the cone. Because of this property, $d_{H}$ is particularly useful for studying the behavior of the rank of the features, which is a scale-invariant function. Indeed, the next result shows that, under mild assumptions, nonnegative weights generate layers that are nonexpansive in Hilbert distance.
@@
Let $A$ be a nonnegative and irreducible matrix with dominant eigenvector $u\in\mathcal{K}$. Assume $X$ to be strictly positive, $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$, and $\sigma$ a continuous (nonlinear) function that is order preserving, subhomogeneous, and such that $u$ is also an eigenvector of $\sigma$. Then
$\max_{i}d_{H}\left(\sigma\big((AXW)_{:,i}\big),u\right)\leq\beta\max_{i}d_{H}(X_{:,i},u),$ (6)
with $0\leq\beta\leq 1$. Where $Y_{:,i}$ denotes the $i$-th column of $Y$. In particular, if $A$ is contractive in the Hilbert distance or $\sigma$ is strictly sub-homogeneous, then $\beta<1$.
In the above statement, order-preserving means that given any $x,y\in\mathcal{K}$ with $x\geq_{\mathcal{K}}y$, it holds $\sigma(x)\geq_{\mathcal{K}}\sigma(y)$, while (strictly) subhomogenenous means that $\sigma(\lambda x)(\ll_{\mathcal{K}})\leq_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and any $\lambda>1$. We recall that, as discussed in Sittoni and Tudisco (2024); piotrowski2024fixed, a broad range of activation functions commonly used in deep learning is subhomogeneous and order-preserving on $\mathcal{K}$. In particular, whenever an activation function is monotone increasing on $\mathbb{R_{+}}$, it is trivially also order-preserving. Additionally, as we prove in Proposition˜A.6, if the activation function is homogeneous, e.g. LeakyReLU, then any nonnegative vector is an eigenvector of $\sigma$. By contrast, if $\sigma$ is strictly subhomogeneous, e.g. $\tanh$, then the only strictly positive eigenvector is the entrywise constant one.
Next, we prove that for neural networks with nonnegative feature representations, the numerical rank goes to 1 as the depth grows to infinity.
@@
Consider a GNN of the form $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(l)}_{:,i}\in\mathcal{K}$ for any $i=1,\dots,d$. If there exists $u\in\mathcal{K}$ such that $\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0$, then
@@
Theorem˜5.3 requires that the Hilbert distance between the feature representation and a fixed vector $u$ goes to zero. Note that this is implied by the relative metrics $E_{\mathrm{Dir}}(X)/\|X\|$ or $E_{\mathrm{Proj}}(X)/\|X\|$ going to zero, but is actually quite weaker. Note also that the bound (6) in Lemma˜5.2 directly provides guidance on situations where the hypotheses of Theorem˜5.3 are satisfied. We discuss several such situations along with some alternative and possibly weaker assumptions for Theorem˜5.3 in detail in Section˜A.6. Finally, we recall that, because of Lemma˜5.2, the last result applies either to GCNs with homogeneous activation function or GATs with any kind of activation function. The convergence of the numerical rank to $1$ may not hold, as discussed in Appendix˜D.
@@
Table 1: Correlation between the classification accuracy and the logarithm of metric values on GNNs with LeakyReLU and depths ranging from 2 to 24 layers, separately trained on different homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings) and large-scale (OGB-Arxiv) datasets. For Erank and NumRank, we subtract 1 so that both metrics approach zero. The rightmost column reports the ratio of classification accuracy between GNNs with 2 and 24 layers. Some heterophilic datasets may be more resilient to the increasing network depth, in-line with observations from the literature, e.g. Guo et al. (2023a). Additional results on other datasets, activation functions and additional network components are presented in Appendix˜E and F.